OpenStudy (anonymous):

Which is a solution to the equation y=3x+1 ? A. (3, 10) B. (2, 6) C. (1, 5) D. (0, 3)

OpenStudy (anonymous):

you can always plug in :) i'm kind of blanking out right now on the more technical way to solve, but if someone can confirm that part, that'd be good :) but you can plug in.. just know that it's like this: (x,y) so sub into the original equation and when one side equals the other, you've got the answer:)

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

no, because that would give you 6=3(2)+1 and 6 is not the same as 7

OpenStudy (anonymous):

I honestly thought i knew how to do this but i don't

OpenStudy (anonymous):

okay.. start with A 10=3(3)+1 10=9+1 10=10 voila!

OpenStudy (anonymous):

So it's A sorry if I'm getting confusing

OpenStudy (anonymous):

haha no it's okay hope that makes sense yeah A

OpenStudy (anonymous):

I was doing it the way you was doing it but it was just confusing to me

OpenStudy (anonymous):

ahh yeah it just takes practice.. just remember (x,y) and plug in respectively :)

OpenStudy (anonymous):

Well thank you for everything

OpenStudy (anonymous):

np:)